how would I go about solving this for example I brought a sofa for $300 and had put 20% down and pay the rest in 8 months How would I find out the amount to pay each month.
Found 2 solutions by nyc_function, richwmiller:
Answer by nyc_function(2741)

$300 times 0.20 = 60
$300 - $60 = $240
Dividing $240 by 8 = $30
The amount to pay each month is $30.

Answer by richwmiller(17219)

.20*300=60 you put 60 down
300-60=240
240/8=30 per month